when the first of two numbers is added to twice the second number, the result is 21, but when the second number is added to twice the first the result is 18. find the two nummbers.: when the first of two numbers is added to twice the second number, the result is 21, but when the second number is added to twice the first the result is 18. find the two nummbers.
Answer by ankor@dixie-net.com(4538)

when the first of two numbers is added to twice the second number, the result is 21, but when the second number is added to twice the first the result is 18. find the two numbers.
:
Let the two numbers be: x & y
:
Write an equation for each statement
"the first of two numbers is added to twice the second number,the result is 21,"
x + 2y = 21
or
x = (21-2y)
:
"the second number is added to twice the first the result is 18."
y + 2x = 18
:
find the two numbers.
:
Substitute (21-2y) for x in the above equation
y + 2(21-2y)= 18
y + 42 - 4y = 18
y - 4y = 18 - 42
-3y = -24
y = (-24)/(-3)
y = +8 is the 2nd number
:
Find x
x = 21 - 2(8)
x = 21 - 16
x = +5
:
Check solutions in the 2nd equation:
y + 2x = 18
8 + 2(5) = 18